OpenStudy (anonymous):

Chad has a bill-fold of one-, five-, and ten-dollar bills totaling $118. He has 2 more than 3 times as many ones as he does five-dollar bills and 1 fewer ten- than five doller bills.How many of each bill does Chad have

OpenStudy (anonymous):

.

hero (hero):

x = # of 1 dollar bills y = # of 5 dollar bills z = # of 10 dollar bills 1x + 5y + 10z = 118 # of one dollar bills is two more than 3 times as many fives x = 3y + 2 # of ten dollar bills is one less than number of fives z = y - 1 3y + 2 + 5y + 10(y - 1) = 118 3y + 5y + 10y + 2 - 10 = 118 18y - 8 = 118 18y = 126 y = 7 So 7 five dollar bills Find the rest

OpenStudy (anonymous):

thank you so much :D

hero (hero):

Did you find the rest?

hero (hero):

You still need to find the # of 1 dollar bills and the # of 10 dollar bills